Of the $50,000 that Natasha pocketed on her last real estate deal, $20,000 went to charity. She invested part of the remainder in Dreyfus New Leaders Fund with an annual yield of 16% and the rest in Templeton Growth Fund with an annual yield of 25%. If she made $6060 on these investments in 1 year, then how much did she invest in each fund?
Answer by checkley77(12844)

50,000-20,000=30,000 for her investments.
.25x+.16(30,000-x)=6,060
.25x+4,800-.16x=6,060
.09x=6,060-4,800
.09x=1,260
x=1,260/.09
x=14,000 invested @ 25%.
30,000-14,000=16,000 invested @ 16%.
Proof:
.25*14,000+.16*16,000=6,060
3,500+2,560=6,060
6,060=6,060